# Enhancing Pre-Trained Generative Language Models with Question Attended Span Extraction on Machine Reading Comprehension

### v2

@@ Abstract
Machine Reading Comprehension (MRC) poses a significant challenge in the field of Natural Language Processing (NLP). While mainstream MRC methods predominantly leverage extractive strategies using encoder-only models such as BERT, generative approaches face the issue of out-of-control generation – a critical problem where answers generated are often incorrect, irrelevant, or unfaithful to the source text. To address these limitations in generative models for extractive MRC, we introduce the Question-Attended Span Extraction (QASE) module. Integrated during the fine-tuning phase of pre-trained generative language models (PLMs), QASE significantly enhances their performance, allowing them to surpass the extractive capabilities of advanced Large Language Models (LLMs) such as GPT-4 in few-shot settings. Notably, these gains in performance do not come with an increase in computational demands. The efficacy of the QASE module has been rigorously tested across various datasets, consistently achieving or even surpassing state-of-the-art (SOTA) results, thereby bridging the gap between generative and extractive models in extractive MRC tasks. Our code is available at this anonymous repo link.
@@
Lin Ai Zheng Hui Zizhou Liu Julia Hirschberg Columbia University, New York, NY {lin.ai, julia}@cs.columbia.edu {zh2483, zl2889@columbia.edu}@columbia.edu
@@ 2 Related Work
Recent MRC research predominantly focuses on extractive question answering using encoder-only PLMs like BERT and XLM-Roberta, predicting the start and end positions of answers directly from the context Ohsugi et al. (2019); Lan et al. (2019); Bachina et al. (2021); Chen et al. (2022). For multi-span answers, Segal et al. (2020) treat this as a sequence tagging task, while others Hu et al. (2019); Lee et al. (2023); Zhang et al. (2023) use hybrid approaches to enhance performance on complex MRC problems. Beyond extractive methods, there is growing interest in applying generative language models for extractive MRC Yang et al. (2020); Li et al. (2021); Jiang et al. (2022); Su et al. (2022), which generate answers by reformulating information across the context.
@@
RAG augments the input of PLMs with in-domain Gu et al. (2018); Weston et al. (2018); Saha and Srihari (2023) or external knowledge Su et al. (2021); Xiao et al. (2021) to control the quality and factual consistency of generated content. It has become a new text generation paradigm in many NLP tasks Li et al. (2022b), such as dialogue response generation Wu et al. (2021); Liu et al. (2023b) and machine translation He et al. (2021); Zhu et al. (2023). However, RAG is typically utilized in scenarios where document retrieval is necessary to reduce input context window Chen et al. (2024); Ram et al. (2023), whereas selective MRC often requires accessing information beyond the immediate context. Our approach diverges from RAG as it directly fine-tunes the weights of the PLMs rather than altering the input to the PLMs with additional information.
@@ 3.1 Question-Attended Span Extraction
To capture the relationship of context tokens to specific questions, we utilize a multi-head attention mechanism (MHA). Each attention head targets different aspects of the context in relation to the question, treating question embeddings as queries and context embeddings as keys and values. Specifically, for each question-context pair, we compute a mean question embedding by averaging the embeddings of question tokens, which is then expanded to align with the length of the context sequence. This expanded question embedding, $z^{*}_{Q}$, serves as the query in the MHA, with the context embedding, $z_{C}$, acting as both key and value. This mechanism allows the derived representation of each token in the context to encapsulate its relevance in relation to the posed question.
@@ 4.3 Does QASE Mitigate Ill-Formed Generation?
Quoref no QASE 45.52 | 52.09 - 58.21 | 63.30 72.77 | 80.90 75.17 | 80.49
(EM | F1) QASE 54.28 | 60.44 - 60.70 | 66.88 75.17 | 81.18 76.19 | 82.13
@@ 4.3.1 QASE-Enhaced PLMs vs SOTA LLMs and Extractive Approaches
On SQuAD, as showed in Table 3, Flan-T5-LargeQASE surpasses human performance, equaling the NLNet model from Microsoft Research Asia and the pre-trained BERT-Large Devlin et al. (2019). Additionally, it surpasses two-shot GPT-4 by 13.6% on EM and 4.0% on F1.
@@ 4.4 Does QASE Improve Factual Consistency?
While token-based EM and F1 scores measure the structural quality of generated text, they do not reflect factual accuracy relative to the context. For this we used $Q^{2}$ Honovich et al. (2021), an automatic metric for assessing factual consistency in generated text, which uses question generation and answering methods over token-based matching. We compared fine-tuned Flan-T5-Large with and without QASE in both single-span (SQuAD) and multi-span (MultiSpanQA) answer settings. Table 6 shows that QASE-enhanced models consistently outperform the vanilla fine-tuned model. On SQuAD, $Q^{2}$ NLI score is improved by 1.0%, and on MultiSpanQA, it is improved by 16.0%.
@@ 4.5 Computational Cost
To assess the computational cost associated with QASE, Table 1 reveals that incorporating the QASE module incurs only a slight increase in the number of trainable parameters in PLMs. The degree of this increase varies based on the hidden sizes of the models. Remarkably, for the largest model, Flan-T5-Large, the addition of QASE accounts for merely an extra 0.2% in parameters. This underscores that QASE can substantially boost the performance of fine-tuned PLMs in MRC tasks without requiring significant additional computational resources.
@@ 4.7 Qualitative Case Studies
Question Attended Alignment
Multi-Span Answers
Flan-T5-LargeQASE also shows a notable improvement in comprehending complex, lengthy sentences and synthesizing answers from information that is sparsely distributed across multiple spans requiring logical processing. This capability is particularly valuable when the answer to a question does not directly stem from a single phrase. Table 9 provides examples of such instances. In Sample 3, the model needs to recognize that ESPN Deportes is the exclusive broadcaster in Spanish and that CBS, although mentioned, does not offer Spanish-language broadcasting. Combining these facts leads to the correct answer, that ESPN Deportes is the network that broadcast the game in Spanish. Flan-T5-LargeQASE accurately generates this answer, whereas Flan-T5-LargeFT incorrectly answers with "CBS", likely due to confusion caused by the complex sentence structures and dispersed information. Similarly, in Sample 4, Flan-T5-LargeQASE correctly identifies the question as seeking the name of the force related to a potential field between two locations. It successfully locates the relevant long sentence, deconstructs, and comprehends it to produce the correct answer, in contrast to Flan-T5-LargeFT, which incorrectly selects the first phrase mentioning "force". In Sample 5, the question asks for the class most commonly not ascribed to the graph isomorphism problem. The model needs to deduce from the context that "it is widely believed that the polynomial hierarchy does not collapse to any finite level", implying "graph isomorphism is not NP-complete". Once again, Flan-T5-LargeQASE arrives at the correct conclusion, while Flan-T5-LargeFT does not.
@@ 5.1 Flan-T5 Zero-Shot Performance
Despite being trained on SQuAD during pre-training, Flan-T5 models demonstrate poor performance across datasets, including SQuAD. While a comprehensive analysis of Flan-T5’s performance is beyond the focus of our study, we briefly explore potential reasons for this underperformance to gain better insights. This underperformance may stem from their training on a wide range of tasks (1,836 tasks), focusing on free-form generation, QA, and reasoning tasks, rather than being finely optimized for extractive QA tasks like MRC. Additionally, generative models like Flan-T5 and Llama 2 generally struggle in MRC tasks, as discussed earlier. For extended discussions, refer to Appendix B.1.
@@ 5.2 Llama 2 Performance
6 Conclusion and Future Work
@@
Due to our limited computational resources, we have been able to perform our experiments on models no larger than Flan-T5-Large. This same constraint led us to only fine-tuning of Llama 2 and Alpaca with LoRA. We note that models based on Llama 2 and Alpaca generally underperform those based on Flan-T5. Apart from the inherent distinctions between decoder-only and encoder-decoder models, and their suitability for different tasks (as seen from the models’ zero-shot performance), a possible factor could be the number of trainable parameters during fine-tuning. Specifically, fine-tuning Llama 2 and Alpaca with LoRA results in only 4.2M trainable parameters, while even the smallest Flan-T5 model provides 77.0M trainable parameters, as shown in Table 1. We acknowledge that many researchers face similar computational resource limitations. Therefore, our research should be very useful, proposing this lightweight module capable of enhancing smaller PLMs to outperform leading LLMs on MRC tasks like these, achieving a balance of effectiveness and affordability.
@@ A.3 Full Experiment Results
Alpaca 15.201 42.759 18.259 33.871 - -
AlpacaFT 52.730 69.099 27.881 43.950 - -
AlpacaQASE 52.196 70.008 37.313 47.622 - -
@@ A.5 Ablation Studies Details
Figure 3 depicts the architecture of the model we use for the ablation studies, with a baseline span extraction module. The baseline span extraction module omits the MHA component, typifying a standard architecture for fine-tuning pre-trained encoders for downstream sequence tagging tasks. The baseline-embedded Flan-T5-Large models are fine-tuned with the same configurations as Flan-T5-LargeQASE including learning rate, weight decay, batch size, epoch number, and GPU type.

### v3

@@ Abstract
Machine Reading Comprehension (MRC) poses a significant challenge in the field of Natural Language Processing (NLP). While mainstream MRC methods predominantly leverage extractive strategies using encoder-only models such as BERT, generative approaches face the issue of out-of-control generation – a critical problem where answers generated are often incorrect, irrelevant, or unfaithful to the source text. To address these limitations in generative models for extractive MRC, we introduce the Question-Attended Span Extraction (QASE) module. Integrated during the fine-tuning phase of pre-trained generative language models (PLMs), QASE significantly enhances their performance, allowing them to surpass the extractive capabilities of advanced Large Language Models (LLMs) such as GPT-4 in few-shot settings. Notably, these gains in performance do not come with an increase in computational demands. The efficacy of the QASE module has been rigorously tested across various datasets, consistently achieving or even surpassing state-of-the-art (SOTA) results, thereby bridging the gap between generative and extractive models in extractive MRC tasks. Our code is available at this GitHub repository.
@@
Lin Ai Zheng Hui Zizhou Liu Julia Hirschberg Columbia University, New York, NY {lin.ai, julia}@cs.columbia.edu {zh2483, zl2889}@columbia.edu
@@ 2 Related Work
Recent MRC research predominantly focuses on extractive question answering using encoder-only PLMs like BERT and XLM-Roberta, predicting the start and end positions of answers directly from the context Ohsugi et al. (2019); Lan et al. (2019); Bachina et al. (2021); Chen et al. (2022). For multi-span answers, Segal et al. (2020) treat this as a sequence tagging task, while others Hu et al. (2019); Lee et al. (2023); Zhang et al. (2023) use hybrid approaches to enhance performance on complex MRC problems. Beyond extractive methods, there is growing interest in applying generative language models for extractive MRC Yang et al. (2020); Li et al. (2021); Jiang et al. (2022); Su et al. (2022), which generate answers by reformulating information across the context. Xu et al. (2021) adopt a similar approach to ours by adding a span extraction auxiliary task to guide text generation. However, this method does not focus sufficiently on the queried questions, which may reduce the accuracy of span extraction.
@@
RAG augments the input of PLMs with in-domain Gu et al. (2018); Weston et al. (2018); Saha and Srihari (2023) or external knowledge Su et al. (2021); Xiao et al. (2021) to control the quality and factual consistency of generated content. It has become a new text generation paradigm in many NLP tasks Li et al. (2022b); Ai et al. (2024), such as dialogue response generation Wu et al. (2021); Liu et al. (2023b) and machine translation He et al. (2021); Zhu et al. (2023). However, RAG is typically utilized in scenarios where document retrieval is necessary to reduce input context window Chen et al. (2024); Ram et al. (2023), whereas selective MRC often requires accessing information beyond the immediate context. Our approach diverges from RAG as it directly fine-tunes the weights of the PLMs rather than altering the input to the PLMs with additional information.
@@ 3.1 Question-Attended Span Extraction
To capture the relationship of context tokens to specific questions, we utilize a multi-head attention mechanism (MHA). Each attention head targets different aspects of the context in relation to the question, treating question embeddings as queries and context embeddings as keys and values. Specifically, for each question-context pair, we compute a mean question embedding by averaging the embeddings of question tokens, which is then expanded to align with the length of the context sequence. This expanded question embedding, $z^{*}_{Q}$, serves as the query in the MHA, with the context embedding, $z_{C}$, acting as both the key and the value. This mechanism allows the derived representation of each token in the context to encapsulate its relevance in relation to the posed question.
@@ 4.3 Does QASE Mitigate Ill-Formed Generation?
Quoref no QASE 45.52 | 52.09 49.05 | 53.81 58.21 | 63.30 72.77 | 80.90 75.17 | 80.49
(EM | F1) QASE 54.28 | 60.44 55.01 | 59.94 60.70 | 66.88 75.17 | 81.18 76.19 | 82.13
@@ 4.3.1 QASE-Enhaced PLMs vs SOTA LLMs and Extractive Approaches
On SQuAD, as showed in Table 3, Flan-T5-LargeQASE surpasses human performance Rajpurkar et al. (2016), equaling the NLNet model from Microsoft Research Asia and the pre-trained BERT-Large Devlin et al. (2019). Additionally, it surpasses two-shot GPT-4 by 13.6% on EM and 4.0% on F1.
@@ 4.4 Does QASE Improve Factual Consistency?
While token-based EM and F1 scores measure the structural quality of generated text, they do not reflect factual accuracy relative to the context. To address this, we used $Q^{2}$ Honovich et al. (2021), an automatic metric for assessing factual consistency in generated text, which uses question generation and answering methods over token-based matching. We compared fine-tuned Flan-T5-Large with and without QASE in both single-span (SQuAD) and multi-span (MultiSpanQA) answer settings. Table 6 shows that QASE-enhanced models consistently outperform the vanilla fine-tuned model. On SQuAD, $Q^{2}$ NLI score is improved by 1.0%, and on MultiSpanQA, it is improved by 16.0%.
@@ 4.5 Computational Cost
To assess the computational cost associated with QASE, Table 1 reveals that incorporating the QASE module incurs only a slight increase in the number of trainable parameters in PLMs. The degree of this increase varies based on the hidden sizes of the models. Remarkably, for the largest model, Flan-T5-Large, the addition of QASE accounts for merely an extra 0.2% in parameters. This underscores the fact that QASE can substantially boost the performance of fine-tuned PLMs in MRC tasks without requiring significant additional computational resources.
@@ 4.7 Qualitative Case Studies
Question Attended Alignment
Multi-Span Answers
Flan-T5-LargeQASE also shows a notable improvement in comprehending complex, lengthy sentences and synthesizing answers from information that is sparsely distributed across multiple spans requiring logical processing. This capability is particularly valuable when the answer to a question does not directly stem from a single phrase. Table 9 provides examples of such instances. In Sample 3, the model needs to recognize that ESPN Deportes is the exclusive broadcaster in Spanish and that CBS, although mentioned, does not offer Spanish-language broadcasting. Combining these facts leads to the correct answer, that ESPN Deportes is the network that broadcast the game in Spanish. Flan-T5-LargeQASE accurately generates this answer, whereas Flan-T5-LargeFT incorrectly answers with “CBS,” likely due to confusion caused by the complex sentence structures and dispersed information. Similarly, in Sample 4, Flan-T5-LargeQASE correctly identifies the question as seeking the name of the force related to a potential field between two locations. It successfully locates the relevant long sentence, deconstructs, and comprehends it to produce the correct answer, in contrast to Flan-T5-LargeFT, which incorrectly selects the first phrase mentioning “force.” In Sample 5, the question asks for the class most commonly not ascribed to the graph isomorphism problem. The model needs to deduce from the context that “it is widely believed that the polynomial hierarchy does not collapse to any finite level,” implying “graph isomorphism is not NP-complete.” Once again, Flan-T5-LargeQASE arrives at the correct conclusion, while Flan-T5-LargeFT does not.
@@
Common Failure Cases
We observe that a recurring error made by Flan-T5-LargeQASE is its inability to correctly interpret Roman numerals, as seen in Failure Sample 1 in Table 11, where the model is asked about the last Super Bowl held in California before Super Bowl 50. The correct answer, “Super Bowl XXXVII,” is clearly mentioned in the context, but Flan-T5-LargeQASE incorrectly identifies “Super Bowl XIX.” The struggle with Roman numeral interpretation leads to errors even when the information is explicitly provided.
Additionally, Flan-T5-LargeQASE sometimes generates slightly redundant phrases, though the excess is minimal. We argue that the generated answers are still correct in the given contexts, and the only reason they are not marked as 100% accurate is due to the dataset’s annotation scheme. For example, in Failure Sample 2 in Table 11, Flan-T5-LargeQASE accurately identifies that the “entrance to studio 5” is the critical element, whereas Flan-T5-LargeFT simplifies the answer to just “studio 5,” missing the nuance. The inclusion of extra words highlights the model’s attention to detail, even though it results in slight deviations from the gold-standard annotations.
We observe minor variations such as the omission or addition of non-essential words like “the” or punctuation marks. While technically considered “mistakes” in strict dataset annotations, these deviations do not affect the semantic accuracy of the model’s responses. These are surface-level discrepancies rather than true comprehension errors. Such differences arguably shouldn’t penalize the model, as they fall within acceptable linguistic flexibility. This prompts a discussion on evaluating model performance, where strict token-level matching may overlook underlying comprehension. For more examples and further discussion, see Appendix A.6.
Failure Sample 1
Context: On May 21, 2013, NFL owners at their spring meetings in Boston voted and awarded the game to Levi’s Stadium. It is the first Super Bowl held in the San Francisco Bay Area since Super Bowl XIX in 1985, and the first in California since Super Bowl XXXVII took place in San Diego in 2003.
Question: Prior to Super Bowl 50, what was the last Super Bowl in California?
Gold Answer: Super Bowl XXXVII
Flan-T5-LargeQASE
Generation
Super Bowl XIX
Flan-T5-LargeFT
Generation
2010
Failure Sample 2
Context: ITV Tyne Tees was based at City Road for over 40 years after its launch in January 1959. In 2005 it moved to a new facility on The Watermark business park next to the MetroCentre in Gateshead. The entrance to studio 5 at the City Road complex gave its name to the 1980s music television programme, The Tube. …
Question: What gave its name to the 1980s music television program “The Tube”?
Gold Answer: The entrance to studio 5
Flan-T5-LargeQASE
Generation
entrance to studio 5 at the City Road complex
Flan-T5-LargeFT
Generation
studio 5
Table 11: Failure cases of Flan-T5-LargeQASE.
@@ 5.1 Flan-T5 Zero-Shot Performance
Despite being trained on SQuAD during pre-training, zero-shot Flan-T5 models demonstrate poor performance across datasets, including SQuAD. While a comprehensive analysis of Flan-T5’s performance is beyond the focus of our study, we briefly explore potential reasons for this underperformance to gain better insights. This underperformance may stem from their training on a wide range of tasks (1,836 tasks), focusing on free-form generation, QA, and reasoning tasks, rather than being finely optimized for extractive QA tasks like MRC. Additionally, generative models like Flan-T5 and Llama 2 generally struggle in MRC tasks, as discussed earlier. For extended discussions, refer to Appendix B.1.
@@ 5.2 Llama 2 Performance
5.3 Performance Discrepancy across Different Base PLMs and Datasets
While we observe significant performance improvements with QASE across various base PLMs and datasets, the extent of the gains varies by dataset. Specifically, Quoref shows the largest improvement, as shown in Table 18 in Appendix B.3, partly due to its weaker baselines. For instance, a Flan-T5-Small model fine-tuned without QASE achieves F1 scores of 85.51% on SQuAD, 76.49% on MultiSpanQA, and 63.30% on Quoref. Higher baseline scores on datasets such as SQuAD make further improvements more challenging but still notable.
Model-wise, Llama 2 and Alpaca show larger improvements than Flan-T5 models, with some exceptions on MultiSpanQA. This is likely due to Flan-T5’s higher baseline performance, as discussed in Sections 5.1, 5.2, and Appendix B.1, B.2. While there are potential factors that explain some of Flan-T5’s superior performance in context-based question answering, a comprehensive comparison of model architectures in MRC tasks is beyond the scope of this study. For further discussion, refer to Appendix B.3.
6 Conclusions and Future Work
@@
Due to our limited computational resources, we have been able to perform our experiments on models no larger than Flan-T5-Large. This same constraint leads us to only fine-tuning of Llama 2 and Alpaca with LoRA. We note that models based on Llama 2 and Alpaca generally underperform those based on Flan-T5. Apart from the inherent distinctions between decoder-only and encoder-decoder models, and their suitability for different tasks (as seen from the models’ zero-shot performance), a possible factor could be the number of trainable parameters during fine-tuning. Specifically, fine-tuning Llama 2 and Alpaca with LoRA results in only 4.2M trainable parameters, while even the smallest Flan-T5 model provides 77.0M trainable parameters, as shown in Table 1. We acknowledge that many researchers face similar computational resource limitations. Therefore, our research should be very useful, proposing this lightweight module capable of enhancing smaller PLMs to outperform leading LLMs on MRC tasks like these, achieving a balance of effectiveness and affordability.
@@ A.3 Full Experiment Results
Alpaca 15.201 42.759 18.259 33.871 9.67 30.02
AlpacaFT 52.730 69.099 27.881 43.950 49.05 53.81
AlpacaQASE 52.196 70.008 37.313 47.622 55.01 59.94
@@ A.5 Ablation Studies Details
Figure 3 depicts the architecture of the model we use for the ablation studies, with a baseline span extraction module. The baseline span extraction module omits the MHA component, typifying a standard architecture for fine-tuning pre-trained encoders for downstream sequence tagging tasks. It closely resembles the approach by Xu et al. (2021), with two key differences: (a) our baseline model integrates both query and context token embeddings to provide additional contextual information, and (b) instead of directly computing the extraction loss, our model includes additional projection and linear layers. The baseline-embedded Flan-T5-Large models are fine-tuned with the same configurations as Flan-T5-LargeQASE including learning rate, weight decay, batch size, epoch number, and GPU type.
@@
A.6 Qualitative Error Analysis
Failure Sample 1
Context: On May 21, 2013, NFL owners at their spring meetings in Boston voted and awarded the game to Levi’s Stadium. It is the first Super Bowl held in the San Francisco Bay Area since Super Bowl XIX in 1985, and the first in California since Super Bowl XXXVII took place in San Diego in 2003.
Question: Prior to Super Bowl 50, what was the last Super Bowl in California?
Gold Answer: Super Bowl XXXVII
Flan-T5-LargeQASE
Generation
Super Bowl XIX
Flan-T5-LargeFT
Generation
2010
Failure Sample 2
Context: The Super Bowl 50 halftime show was headlined by the British rock group Coldplay with special guest performers Beyonce and Bruno Mars, who headlined the Super Bowl XLVII and Super Bowl XLVIII halftime shows, respectively.
Question: At which Super Bowl did Beyonce headline the halftime show?
Gold Answer: Super Bowl XLVII
Flan-T5-LargeQASE
Generation
Super Bowl 50
Flan-T5-LargeFT
Generation
Super Bowl 50
Table 15: Failure cases of Flan-T5-LargeQASE in interpreting Roman numerals.
Challenges in Roman Numeral Interpretation
We observe that a recurring error made by Flan-T5-LargeQASE is its inability to correctly interpret Roman numerals, as evidenced in Table 15. In Failure Sample 1, the model is asked about the last Super Bowl held in California before Super Bowl 50. The correct answer, “Super Bowl XXXVII,” is clearly mentioned in the context, but Flan-T5-LargeQASE incorrectly identifies “Super Bowl XIX.” Similarly, in Failure Sample 2, the context states that Beyonce headlined Super Bowl XLVII, yet the model incorrectly identifies “Super Bowl 50” as the answer, despite the clear mention of Super Bowl XLVII in the question and context. These examples indicate that Flan-T5-LargeQASE struggles with Roman numeral interpretation, leading to errors even when the information is explicitly provided in the text.
Failure Sample 3
Context: Super Bowl 50 was an American football game to determine the champion of the National Football League (NFL) for the 2015 season. The game was played on February 7, 2016, at Levi’s Stadium in the San Francisco Bay Area at Santa Clara, California. …
Question: What city did Super Bowl 50 take place in?
Gold Answer: Santa Clara
Flan-T5-LargeQASE
Generation
Santa Clara, California
Flan-T5-LargeFT
Generation
San Francisco Bay Area
Failure Sample 4
Context: ITV Tyne Tees was based at City Road for over 40 years after its launch in January 1959. In 2005 it moved to a new facility on The Watermark business park next to the MetroCentre in Gateshead. The entrance to studio 5 at the City Road complex gave its name to the 1980s music television programme, The Tube. …
Question: What gave its name to the 1980s music television program “The Tube”?
Gold Answer: The entrance to studio 5
Flan-T5-LargeQASE
Generation
entrance to studio 5 at the City Road complex
Flan-T5-LargeFT
Generation
studio 5
Table 16: Failure cases of Flan-T5-LargeQASE in generating redundant phrases.
Minor Redundancies in Generation
Another common error we observe is that Flan-T5-LargeQASE tends to generate slightly redundant phrases, though the excess is minimal. We argue that the generated answers are still correct in the given contexts, and the only reason they are not marked as 100% accurate is due to the dataset’s annotation scheme. In Failure Sample 3, for instance, Flan-T5-LargeFT produces a completely incorrect answer, while Flan-T5-LargeQASE provides the correct answer, but with a minor additional word, “California,” which does not detract from its correctness. Similarly, in Failure Sample 4, Flan-T5-LargeQASE accurately identifies that the “entrance to studio 5” is the critical element, whereas Flan-T5-LargeFT simplifies the answer to just “studio 5,” missing the nuance. The inclusion of extra words in Flan-T5-LargeQASE’s responses highlights the model’s attention to detail, even though it results in slight deviations from the gold-standard annotations, as shown in Table 16.
Surface-Level Variations in Output
Another pattern we observe involves minor variations in Flan-T5-LargeQASE’s outputs, such as the omission or addition of non-essential words like “the” or punctuation marks, such as quotation marks. These deviations, while technically considered “mistakes” in the strict context of dataset annotations, do not alter the semantic accuracy of the model’s responses. For example, if the model omits a definite article or does not replicate quotation marks around a phrase, the intended meaning of the answer remains intact. These variations are more reflective of surface-level discrepancies rather than true comprehension errors. One could argue that such differences should not penalize the model, as they fall within the bounds of acceptable linguistic flexibility. This raises an interesting discussion about how we evaluate model performance, particularly in cases where strict adherence to token-level matching might overlook the model’s underlying comprehension. These “errors” suggest that Flan-T5-LargeQASE generates answers with a focus on meaning rather than perfect alignment with rigid annotation structures, thus offering a more human-like adaptability in its output.